OpenStudy (anonymous):

Decide whether or not the order pair (6,-2) is a solution of the system. 3x=-20-y 2x=-18-3y

OpenStudy (anonymous):

3(6)=-20-(-2) 18=-18 Not correct, not a solution 2(6)=-18-3(-2) 12=-12 Not a solution